If I fill 3 molecules of air which are very hot in a container and if I fill 10 molecules of air which have room temp in the another container which is exactly identical to first one that is has same volume so which container is gonna experience more pressure on its walls
The one which has high temp 3 molecules
Or the other having room temp 10 molecules
Actually 3 molecules have been heated while those 10 have not been
1
Expert's answer
2019-05-08T09:30:35-0400

Of course, it does not make sense to talk about pressure, which surrounds 3 or 10 or even 1000 molecules, there should be a lot more.

Nevertheless, since pressure p = nkT, therefore

"\frac{p_1}{p_2} = \frac{N_1T_1}{N_2T_r}"

and if

"T_1>\frac{T_r10}{3}"

respectively

"p_1>p_2"